1. A man in a boat starts rowing; and it accelerates from 0 to 0.80 m/s in a distance of 0.51m. If the combined mass of the person and the boat is 85kg, what is the magnitude of the net force acting on the boat?
1
Expert's answer
2021-04-01T18:33:20-0400

Let's first find the acceleration of the boat:


"v^2=v_0^2+2as,""a=\dfrac{v^2-v_0^2}{2s},""a=\dfrac{(0.8\ \dfrac{m}{s})^2}{2\cdot0.51\ m}=0.63\ \dfrac{m}{s^2}."

Finally, we can find the net force acting on the boat:


"F_{net}=ma=85\ kg\cdot0.63\ \dfrac{m}{s^2}=53.55\ N."